p,q are the roots of  mx.x+x(2-m)+3. g,h are two values of m satisfying the equation p/q +q/p =2/3 then the value of g/hh  +h/gg 

Grade:

1 Answers

Roshan Mohanty
10 years ago

we have p^2 + q^2/(pq)=2/3
=>(p+q)^2 - 2pq /(pq)=2/3
p+q=m-2
pq=3
so we get (m-2)^2 = 8
so m= 2 + root(8)[g] or 2-root(8)[h]
then g/h + h/g = -4